OpenStudy (anonymous):

Please help me I need help with a few homework questions, I will medal and fan!

OpenStudy (anonymous):

Bonnie has a container in the shape of a rectangular pyramid. The formula for the surface area of the enclosed space is S = lw + 0.5Ph. Solve for P. P = S – lw – 0.5h P = S + lw + 0.5h P = The quantity S minus l times w all divided by 0.5 times h P = S divided by the quantity l times w plus 0.5 times h

OpenStudy (sachintha):

First take Iw to the left hand side.

OpenStudy (sachintha):

Give it a try.

OpenStudy (anonymous):

i have tried, but I suck :( @Sachintha do you think you could walk me through it?

OpenStudy (sachintha):

For the first step you remove Iw by the right hand side by subtracting Iw from both sides. \[S - Iw = Iw+0.5Ph-Iw\]

OpenStudy (sachintha):

So Iw on right hand side cancel out.

OpenStudy (anonymous):

ok

OpenStudy (sachintha):

Then you are left with, \[S-Iw=0.5Ph\]

OpenStudy (anonymous):

ok. next would you add Iw to the other side? to cancel out?

OpenStudy (sachintha):

No. What we are trying to do is to get an equation for P. So P is left at right hand side.

OpenStudy (anonymous):

O.K!?

OpenStudy (sachintha):

Finally you divide both sides by 0.5h. \[\frac{S-Iw}{0.5h}=\frac{0.5Ph}{0.5h}\] \[P=\frac{S-Iw}{0.5h}\]

OpenStudy (anonymous):

so c?

OpenStudy (sachintha):

Yeah you are correct. :)

OpenStudy (sachintha):

Hope you understood every step.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

If I post another question, will you help with it?

OpenStudy (anonymous):

I will open a new question so I can give you another medal

OpenStudy (anonymous):

oh wow lol